OpenStudy (anonymous):

3⁻³ x 3⁴ x 3 x 3⁻⁵ A.) –27 B.) 1/27 C.) –81 D.) 1/81

OpenStudy (anonymous):

3^-3=1/3*3*3

OpenStudy (anonymous):

b

OpenStudy (anonymous):

3^4=3*3*3*3

OpenStudy (anonymous):

Thank you !

OpenStudy (anonymous):

you're welcome :)